OpenStudy (anonymous):

Which value of b from the set {1, 2, 3, 4, 5} makes this equation true? 2b + 3(b – 3) = 6

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

yes you are right because 2b+3(b−3)=6 Step 1: Simplify both sides of the equation. 2b+3(b−3)=6 Simplify: (Show steps) 5b−9=6 Step 2: Add 9 to both sides. 5b−9+9=6+9 5b=15 Step 3: Divide both sides by 5. 5b5=155 b=3 Answer: b=3

OpenStudy (anonymous):

:)